Name two points on a line that has a slope of 5/8.
Answer by macston(5194)

slope=m=5/8
y=mx+b
y=(5/8)x+b Pick any point (I'll pick (0,0))
0=(5/8)0+b
0=b
For our line, (0,0) is one point and b=0
y=5/8x Pick a value for x(I'll pick x=8)
y=(5/8)(8)
y=5
Our second point is (8,5)
.
ANSWER: Two points on a line with slope 5/8: (0,0),(8,5)